Mathematics 242 – Principles of Analysis Definitions Quiz 4 – March 4, 2011

Your Name:

A) (10) What is the definition of the property that a sequence {x_{n}} converges to a ∈ R?

B) (10) What is the definition of the least upper bound of a nonempty bounded set A ⊂ R?